OpenStudy (anonymous):

I am having problems trying to figure out how to solve radical expressions theproblem is 3(checkmarkthing) x-2=3

OpenStudy (anonymous):



OpenStudy (anonymous):

Don't you solve for x?

OpenStudy (anonymous):

yes but i cant figure out how to get x alone im lookint at my book and i dont understand it

OpenStudy (anonymous):

ah ok, well you know your order of operations right? when you are simplifying, all you do is instead of PEMDAS you SADMEP

OpenStudy (anonymous):

so add/subtract from the x side, then multiply/divide then square both sides

OpenStudy (anonymous):



OpenStudy (anonymous):

erm, no 3(sqrtx)-2=3 +2 +2 3(sqrt x)=5 /3 /3 (sqrt x)=5/3 and in order to get rid of a square root you square it so ou would get (-/x)^2=(5/3)^2

OpenStudy (anonymous):

is it \[3\sqrt{x-2}=3\] or is it \[3\sqrt{x}-2=3\]?

OpenStudy (anonymous):

I think the second one, juding by the pic at the top

OpenStudy (anonymous):

firstone

OpenStudy (anonymous):

oh ok then, the you get 3(sqrt x-2)=2 /3 /3

OpenStudy (anonymous):

and square both sides

OpenStudy (anonymous):

then add 2

OpenStudy (anonymous):

then divide both sides by 3 and get \[\sqrt{x-2}=1\] square both sides and get \[x-2=1\] therefor \(x=3\)

OpenStudy (anonymous):

^3sqrt x-2=3